OpenStudy (oswaldmurphy):

Given the function g(x) = 8x − 2, compare and contrast g(−2) and g(4). Choose the statement that is true concerning these two values. The value of g(−2) is larger than the value of g(4). The value of g(−2) is the same as the value of g(4). The value of g(−2) is smaller than the value of g(4). The values of g(−2) and g(4) cannot be compared.

OpenStudy (oswaldmurphy):

@Mehek14

jimthompson5910 (jim_thompson5910):

To find the value of g(-2) you replace every copy of x with -2. Then you simplify

jimthompson5910 (jim_thompson5910):

g(x) = 8x - 2 g(-2) = 8(-2) - 2 ... replace every x with -2 g(-2) = -16 - 2 g(-2) = -18 hopefully this makes sense?

OpenStudy (oswaldmurphy):

Why -2, because that is the first one of the two so it = x?

jimthompson5910 (jim_thompson5910):

because it says "contrast g(-2) and g(4)"

jimthompson5910 (jim_thompson5910):

I showed how to get g(-2), which is -18 are you able to find g(4) ?

OpenStudy (oswaldmurphy):

Oh so I have to do both. Ok it makes since now. g(4) = 30?

jimthompson5910 (jim_thompson5910):

g(4) = 30 is correct

OpenStudy (oswaldmurphy):

So the answer is C?

jimthompson5910 (jim_thompson5910):

yep ` The value of g(−2) is smaller than the value of g(4).` is correct

OpenStudy (oswaldmurphy):

Thanks!

jimthompson5910 (jim_thompson5910):

no problem